OpenStudy (taylor):

x-y=5 can anyone solve this? or explain how to solve this problem?

OpenStudy (anonymous):

I believe you need to solve for y to then continue to solve for x

OpenStudy (anonymous):

no...thats not right...

OpenStudy (anonymous):

if you have an equation with 2 unknown variables, you need 2 different equations to solve it...

OpenStudy (anonymous):

the equation you gave is for a graph...there are a lot of possible questions that can come from graphs...such as what is the x or y intercept

OpenStudy (taylor):

im only solving for y

OpenStudy (sandra):

right, so to solve for y, that means you want to get y alone on one side of the equation

OpenStudy (oaktree):

I guess you could say that y=x-5, but past that you can't do anything without another equation.

OpenStudy (anonymous):

y=x-5...this means that the y intercept on a graph is -5 and the x intercept is 5...the slope is 1...etc

OpenStudy (anonymous):

hope that helps.

OpenStudy (sandra):

right so the basic rule of algebra is, you can do the same operations to both sides o the equation and know that it still holds - this means you can add anything to both sides, subtract, divide, or multiple both sides by the same thing and it's still true

OpenStudy (sandra):

so x-y = 5

OpenStudy (sandra):

so get y alone, you could try adding "y" o both sides

OpenStudy (anonymous):

yes and subtracting 5 from both sides as well

OpenStudy (sandra):

so 1. x - y = 5 2. x - y + y = 5 + y 3. x = 5 + y 4. and no subtract y from both sides 5. x - 5 = 5 + y -5 6. x - 5 = y 7. y = x -5

OpenStudy (taylor):

ok I think I got it now...thanks so much

OpenStudy (anonymous):

no=now* ;)

OpenStudy (anonymous):

and your subtracting 5..not y :p

OpenStudy (sandra):

lol yep!

OpenStudy (sandra):

ack!